How do i write the equation of a line,in slope-intercept form, with slope 1/2 and y-intercept -9.
Found 2 solutions by checkley71, ilana:
Answer by checkley71(8403)

Y=mX+b
Y=X/2-9 ANSWER.

Answer by ilana(307)

Slope-intercept form is y=mx+b, where m is the slope and b is the y-intercept. So you plug in your m and b and get y=(1/2)x+(-9), or y=(1/2)x-9.